# Simulation as Supervision: Mechanistic Pretraining for Scientific Discovery

### v3

@@ 1 Abstract
Scientific modeling faces a tradeoff between the interpretability of mechanistic theory and the predictive power of machine learning. While hybrid approaches like Physics-Informed Neural Networks (PINNs) embed domain knowledge as functional constraints, they can be brittle under model misspecification. We introduce Simulation-Grounded Neural Networks (SGNNs), a framework that instead embeds domain knowledge into the training data to establish a structural prior.
By pretraining on synthetic corpora spanning diverse model structures and observational artifacts, SGNNs learn the broad patterns of physical possibility. This allows the model to internalize the underlying dynamics of a system without being forced to satisfy a single, potentially incorrect equation.
We evaluated SGNNs across scientific disciplines and found that this approach confers significant robustness. In prediction tasks, SGNNs nearly tripled COVID-19 forecasting skill versus CDC baselines. In tests on dengue outbreaks, SGNNs outperformed physics-constrained models even when both were restricted to incorrect human-to-human transmission equations, demonstrating that SGNNs are potentially more robust to model misspecification. For inference, SGNNs extend the logic of simulation-based inference to enable supervised learning for unobservable targets, estimating early COVID-19 transmissibility more accurately than traditional methods. Finally, SGNNs enable back-to-simulation attribution, a form of mechanistic interpretability that maps real-world data back to the simulated manifold to identify underlying processes. By unifying these disparate simulation-based techniques into a single framework, we demonstrate that mechanistic simulations can serve as effective training data for robust scientific inference that generalizes beyond the limitations of fixed functional forms.
@@ 2 Introduction
Scientific modeling is often divided between two approaches: theory-driven models, which encode domain knowledge into mathematical equations, and data-driven models, which learn flexible patterns from large datasets. Each has fundamental limitations. Mechanistic models are interpretable and grounded in scientific theory, but often break down in practice due to noise, unobserved variables, or incorrect structural assumptions. Neural networks can capture complex relationships but require large labeled datasets, lack scientific grounding, and often learn spurious patterns that fail when applied to new data. We lack models that are both scientifically grounded and capable of robust generalization from imperfect data.
@@
Recent work has sought to bridge this divide by incorporating mechanistic knowledge into machine learning. These approaches can be understood through the three core components of machine learning problems: the data, the model architecture, and the optimization procedure. Physics-Informed Neural Networks (PINNs) incorporate governing equations into the loss function, constraining optimization to satisfy known physical laws [1, 2]. Neural ODEs embed continuous-time dynamics directly into the architecture, parameterizing derivative functions that are solved during forward passes [3]. Universal Differential Equations (UDEs) combine both approaches, using neural networks to learn unknown terms in otherwise known equations while training on trajectory data [4]. These methods act as functional constraints: they perform well when equations are known and all relevant variables are observable, but can introduce bias or cause the learning process to fail when mechanistic models are incomplete or structurally misspecified.
Another approach is to incorporate mechanistic knowledge into the training data itself: using simulations from mechanistic models to generate synthetic datasets for neural network training. This strategy has emerged across disciplines, driven by different motivations. In physics, simulation-trained networks can serve as fast surrogates for expensive simulators, enabling real-time inference where full simulations are computationally prohibitive [5]. In epidemiology and social science, simulations address data scarcity and enable learning of unobservable quantities [6]. Despite strong empirical results in specific applications, these efforts lack a unified framework for understanding when and why simulation-grounded learning succeeds.
@@
We introduce Simulation-Grounded Neural Networks (SGNNs) as a general framework for training neural networks on mechanistic simulations. While simulation-grounded learning has been applied in isolated contexts, we provide a unified treatment and evaluation across domains. Unlike approaches that enforce pointwise constraints, SGNNs embed domain knowledge into the training data to establish a structural prior. By pretraining on synthetic corpora spanning diverse model structures, parameter regimes, and realistic observational artifacts, the model internalizes the global geometry of physically plausible dynamics. This allows the neural network to inherit scientific structure from the underlying math without being explicitly constrained by a single, potentially incorrect governing equation. This resulting framework is architecture-agnostic and maintains flexibility across diverse scientific domains.
SGNNs are aligned with recent work demonstrating that large neural networks can acquire general-purpose inference strategies through synthetic supervision [7, 8, 9]. These models show that abstract generative processes can produce strong reasoning priors. SGNNs build on this foundation but differ fundamentally in being grounded in domain-specific scientific simulations. This yields models that are not only performant, but mechanistically interpretable and aligned with scientific theory. In this sense, SGNNs act as scientific inference engines: trained on simulations, and capable of reasoning under real-world uncertainty.
@@
This approach directly addresses a core limitation of machine learning in science: the inability to infer unobservable quantities. Scientific parameters like the basic reproduction number in an infectious disease outbreak or ecological carrying capacity can’t be measured directly. Traditional methods estimate them by fitting mechanistic models to observed trajectories, a process sensitive to noise, delays, and structural mismatch. Simulation-based inference (SBI) addresses this by using simulation-grounded training to learn posterior distributions over parameters [10, 11, 12], providing a principled approach to Bayesian parameter inference. However, full posterior inference is not always necessary: simulation-grounded learning can also enable simpler approaches such as direct quantile regression on parameters, which can achieve accurate inference with reduced computational complexity.
@@
We demonstrate SGNNs across tasks in epidemiology, ecology, chemistry, and social science, showing strong generalization in forecasting, regression, classification, and parameter inference. SGNNs nearly triple the median forecasting skill of the CDC Forecast Hub on early COVID-19 data, despite never seeing real-world disease data. Additionally, SGNNs trained only on human-to-human disease simulations outperform specialized models on dengue (a vector-borne disease) by 24%, demonstrating the robustness of the approach. In chemical reaction yield prediction, SGNNs reduce residual variance by one-third, and in ecological forecasting, SGNNs retain high accuracy in high-dimensional settings where task-specific neural nets fail completely.
These results characterize SGNNs as a general framework for scientific machine learning, demonstrating when and how mechanistic simulations can serve as effective training data for robust, interpretable inference across disciplines. [^1] Code and data available at https://github.com/carsondudley1/SGNNs.
3 The Simulation-Grounded Neural Network (SGNN) Framework
@@ 3.1 Conceptual Framework
SGNNs train neural networks on synthetic data generated by mechanistic simulators. Let $S(\theta)$ denote a stochastic simulator parameterized by $\theta$, which generates trajectories or time series $x_{i}=S(\theta_{i})$. To capture a wide range of system behaviors, we sample $\theta_{i}\sim P(\theta)$ from a task-specific distribution that spans multiple model classes, parameter regimes, and observational settings.
Each simulation yields one or more training examples $x_{i}$ paired with targets $y_{i}=T(\theta_{i})$, where $T(\cdot)$ defines the labeling function (e.g., future trajectories, final states, or latent scientific quantities). Importantly, simulations are passed through an observation model that adds realistic artifacts such as noise, censoring, delays, or reporting bias. The result is a synthetic dataset $\mathcal{D}_{\text{syn}}=\{(x_{i},y_{i})\}$ that reflects both mechanistic diversity and realistic surveillance scenarios.
A neural network $f_{\phi}$ is then trained via supervised learning to predict $y_{i}$ from $x_{i}$. Because each input is paired with exact labels, including latent scientific quantities, and the training set spans a wide range of mechanistic regimes and observational conditions, the model learns to internalize structure that generalizes beyond any single model or parameter setting. SGNNs are architecture-agnostic: any neural network with sufficient capacity may serve as $f_{\phi}$. In practice, we use domain-specific backbones such as temporal transformers, graph networks, or hybrid models (Appendix B).
This simulation-grounded approach supports direct supervision for both observable and unobservable targets, and enables out-of-the-box generalization to real-world data.
@@ 3.2 Key Capabilities
Simulation-Grounded Neural Networks (SGNNs) address core limitations of both mechanistic and data-driven modeling approaches. The framework offers five key capabilities:
@@
  • Robustness to misspecification. By training on an ensemble of models spanning multiple structures, parameter regimes, and noise processes, SGNNs develop flexible representations that generalize across regimes—even when real-world dynamics deviate from any one simulator.
@@
  • Mechanistic interpretability. Through back-to-simulation attribution, SGNNs identify which simulated regimes they internally associate with real-world inputs, offering insight into the underlying dynamics the model believes are active, and enabling process-level explanations.
@@ 4.1 Overview of Evaluation Strategy
We evaluate Simulation-Grounded Neural Networks (SGNNs) across forecasting, inference, regression, and classification tasks spanning multiple scientific domains. Our experiments are designed to assess four core capabilities: (i) generalization to real-world prediction under limited or noisy data, (ii) supervised learning for unobservable scientific parameters via synthetic supervision, (iii) applicability across domains and task types, and (iv) the role of mechanistic grounding in driving model performance.
4.2 SGNNs Enable Robust Infectious Disease Forecasting Across Regimes
@@ 4.2.1 SGNNs Outperform CDC COVID-19 Forecasting Models
We assessed performance using forecasting skill, defined as the percentage improvement in mean absolute error (MAE) relative to a naive baseline that predicts the most recent value:
@@
where the baseline prediction is $\hat{y}_{t+h}=y_{t}$ for all forecast horizons $h$. This is a stringent metric in infectious disease forecasting. The mean skill across all models in the CDC Forecast Hub was negative (–1.0) [16]. For comparison, we benchmarked the SGNN against both the median skill (13.0) and best model’s skill (34.0) of Forecast Hub models.
The SGNN outperformed the baselines (Figure 2). Compared to Chronos [17], a state-of-the-art foundation model pretrained on generic time series, the SGNN achieved a roughly 12-fold improvement in skill, underscoring the importance of mechanistic grounding. The SGNN also nearly tripled the Forecast Hub’s median skill and slightly exceeded its best model performance. Finally, the SGNN outperformed a physics-informed neural network (PINN) baseline based on SEIR equations (skill = –112), which performed poorly in our tests in the presence of noise and data sparsity, consistent with prior observations [2]. This demonstrates that incorporating domain knowledge into the data rather than constraining the output of a neural network to match specific equations may be a better strategy in the presence of large uncertainty and minimal real-world data.
Figure 2: SGNNs outperform state-of-the-art baselines in real-world disease forecasting and reveal the importance of mechanistic grounding. (A) SGNNs achieve 35.3% forecasting skill on early COVID-19 mortality, almost tripling the CDC Forecast Hub median and exceeding its best model—despite using no real COVID-19 data. (B) SGNNs produce accurate forecasts with calibrated uncertainty across diverse real-world locations, including Michigan, New York, Texas, and California. (C) Across all U.S. states, SGNNs consistently achieve lower error (lighter) than Chronos and PINNs, showing robust generalization from synthetic pretraining. (D) On dengue, a domain with fundamentally different transmission dynamics, SGNNs outperform both PINNs and statistical models in a fully zero-shot setting—demonstrating robustness to mechanistic misspecification. (E) Replacing mechanistic simulations with neural simulators causes performance collapse, highlighting the necessity of mechanistic fidelity. (F) Restricting pretraining to simple SEIR models leads to overconfident exponential forecasts, confirming that mechanistic diversity and surveillance realism are essential for robust downstream performance.
@@ 4.2.2 SGNNs Forecast Accurately Despite Mechanistic Misspecification
Mechanistic models—whether traditional compartmental models or hybrid approaches like PINNs—rely on structural assumptions that may not hold in practice. Real-world systems often deviate from these assumptions, leading to brittle forecasts when the model form is even slightly wrong. This challenge is especially acute when key mechanisms (e.g., transmission routes or feedback loops) are misrepresented or omitted. To assess SGNNs under such conditions, we tested their forecasting performance with different mechanisms from those seen during pretraining.
@@
These findings underscore a strength of SGNNs: by training on a broad ensemble of mechanistic simulations spanning multiple model structures and parameter regimes, SGNNs learn flexible, scientifically grounded representations that can generalize when structural elements differ from those seen in training. This result suggests that SGNNs do not merely memorize surface patterns, but instead internalize scientific principles that extend across mechanistic boundaries. SGNNs may be learning abstract representations that capture core features of disease spread—such as delays, feedback, and depletion—that hold across different mechanistic forms.

### v4

@@ 1 Abstract
Scientific modeling faces a tradeoff between the interpretability of mechanistic theory and the predictive power of machine learning. While existing hybrid approaches have made progress by incorporating domain knowledge into machine learning methods as functional constraints, they can be limited by a reliance on precise mathematical specifications. When the underlying equations are partially unknown or misspecified, enforcing rigid constraints can introduce bias and hinder a model’s ability to learn from data. We introduce Simulation-Grounded Neural Networks (SGNNs), a framework that incorporates scientific theory by using mechanistic simulations as training data for neural networks. By pretraining on diverse synthetic corpora that span multiple model structures and realistic observational noise, SGNNs internalize the underlying dynamics of a system as a structural prior.
We evaluated SGNNs across multiple disciplines, including epidemiology, ecology, social science, and chemistry. In forecasting tasks, SGNNs outperformed both standard data-driven baselines and physics-constrained hybrid models. They nearly tripled the forecasting skill of the average CDC models in COVID-19 mortality forecasts and accurately forecasted high-dimensional ecological systems. SGNNs demonstrated robustness to model misspecification, performing well even when trained on data with incorrect assumptions. Our framework also introduces back-to-simulation attribution, a method for mechanistic interpretability that explains real-world dynamics by identifying their most similar counterparts within the simulated corpus. By unifying these techniques into a single framework, we demonstrate that diverse mechanistic simulations can serve as effective training data for robust scientific inference.
@@ 2 Introduction
Scientific modeling is often divided between two approaches: theory-driven models, which encode domain knowledge into mathematical equations, and data-driven models, which learn flexible patterns from large datasets. Each has fundamental limitations. Mechanistic models are interpretable and grounded in scientific theory, but often break down in practice due to noise, unobserved variables, or incorrect structural assumptions. Neural networks can capture complex relationships but require large labeled datasets, lack scientific grounding, and often learn spurious patterns that fail when applied to new data. There remains a critical need for modeling frameworks that are both scientifically grounded and capable of robust generalization from imperfect real-world data.
@@
Recent work has sought to bridge this divide by incorporating mechanistic knowledge into machine learning. These approaches can be understood through the three core components of machine learning problems: the data, the model architecture, and the optimization procedure. Physics-Informed Neural Networks (PINNs) incorporate governing equations into the loss function, constraining optimization to satisfy known physical laws [34, 36]. Universal Differential Equations (UDEs) incorporate domain knowledge into the model architecture by using neural networks to represent unknown terms within a larger system of differential equations [33]. These methods perform well when equations are known and all relevant variables are observable, but can introduce bias or cause the learning process to fail when mechanistic models are incomplete or structurally misspecified.
Another approach is to incorporate mechanistic knowledge into the training data itself by using simulations to generate synthetic datasets for neural network training. This strategy has emerged across several disciplines: in physics, simulation-trained networks often serve as fast surrogates for computationally expensive models [30, 2], while in epidemiology and social science, they can address data scarcity issues by generating synthetic trianing data [46]. However, many of these efforts rely on a single, fixed simulator, making them brittle in the same sense as methods that directly constrain neural network outputs. While training on a single model structure is effective in domains like physics where governing equations are well-understood, it is less effective in fields where the data-generating process is fundamentally uncertain. In an emerging pandemic, for instance, a model restricted to a single mechanistic hypothesis cannot account for the vast range of plausible transmission dynamics, causing the neural network to memorize the specific biases of that model. Despite promising empirical results in isolated contexts, these approaches lack a unified framework for achieving robustness through mechanistic diversity.
@@
We introduce Simulation-Grounded Neural Networks (SGNNs) as a general framework for training neural networks on ensembles of mechanistic simulations. While simulation-grounded learning has been applied in a range of different contexts, our goal here is to provide a unified treatment and evaluation across domains. By pretraining on synthetic corpora spanning diverse model structures, parameter regimes, and realistic observational artifacts, the model internalizes the global geometry of physically plausible dynamics. This allows the neural network to inherit scientific structure from the underlying models without being explicitly constrained by a single, potentially misspecified governing model. Our framework is architecture-agnostic and maintains flexibility across diverse scientific domains.
SGNNs are aligned with recent work demonstrating that large neural networks can acquire general-purpose inference strategies through training on synthetic data [27, 18, 39]. However, while previous models in this space rely on generic statistical priors to generate training data—such as Gaussian processes or broad distributions of time series—SGNNs use domain-specific scientific simulations to generate a training corpus. This shift from statistical to mechanistic supervision ensures that the model internalizes the causal structure and constraints of the physical world. Consequently, SGNNs yield models that are not only performant but also mechanistically interpretable and inherently aligned with scientific theory.
@@
This approach directly addresses a core limitation of machine learning in science: the inability to infer unobservable quantities. Scientific parameters like the basic reproduction number in an infectious disease outbreak or ecological carrying capacity cannot be measured directly. Traditional methods estimate them by fitting mechanistic models to observed trajectories, a process sensitive to noise, delays, and structural mismatch. Simulation-based inference (SBI) [14, 22, 9] addresses this by using simulation-grounded training to learn posterior distributions over parameters, providing a principled approach to Bayesian parameter inference. However, full posterior inference is not always necessary: simulation-grounded learning can also enable simpler approaches such as direct quantile regression on parameters, which can achieve accurate inference with reduced computational complexity.
@@
We evaluate the SGNN framework across four distinct scientific domains: epidemiology, ecology, chemistry, and social science. Our experiments cover a broad range of tasks, including time-series forecasting, parameter inference, yield regression, and network classification. In ecological forecasting, SGNNs maintain robust accuracy in relatively high-dimensional settings (up to 16 species) where task-specific neural networks often fail to generalize. In chemical synthesis, the model reduces residual variance in reaction yield prediction by one-third compared to standard data-driven baselines. For infectious disease forecasting, SGNNs pretrained on synthetic data demonstrate competitive performance against real-world benchmarks, nearly tripling the median skill of the CDC Forecast Hub during the early COVID-19 pandemic and outperforming specialized models on vector-borne dengue outbreaks by 24%. Finally, in social network analysis, SGNNs accurately identify the origin of diffusion cascades under conditions of partial observability. Together, these results suggest that mechanistic simulations can provide a robust foundation for inference across disparate scientific tasks.  [^1] Code and data available at https://github.com/carsondudley1/SGNNs.
3 Simulation-Grounded Neural Network (SGNN) Framework
@@ 3.1 Conceptual Framework
SGNNs train neural networks on synthetic data generated by mechanistic simulators. Let $S(\theta)$ denote a stochastic simulator parameterized by $\theta$, which generates trajectories or time series $x_{i}=S(\theta_{i})$. $\theta$ sets the parameter values and decides which of multiple mechanisms will be turned on for each simulation. To capture a wide range of system behaviors, we sample $\theta_{i}\sim P(\theta)$ from a task-specific distribution that spans multiple model classes, parameter regimes, and observational settings. Importantly, simulations are passed through an observation model that adds realistic artifacts such as noise, censoring, delays, or reporting bias.
Each simulation yields one or more training examples $x_{i}$ paired with targets $y_{i}=T(\theta_{i})$, where $T(\cdot)$ defines the labeling function (e.g., future trajectories, final states, or latent scientific quantities). The result is a synthetic dataset $\mathcal{D}_{\text{syn}}=\{(x_{i},y_{i})\}$ that reflects both mechanistic diversity and realistic surveillance scenarios.
A neural network $f_{\phi}$ is then trained via supervised learning to predict $y_{i}$ from $x_{i}$. Because each input is paired with exact labels, including latent scientific quantities, and the training set spans a wide range of mechanistic regimes and observational conditions, our aim is that the model learns to internalize structure that generalizes beyond any single model or parameter setting. SGNNs can be built using a range of possible architectures, so long as they have sufficient capacity. In practice, we use domain-specific backbones such as temporal transformers, graph networks, or hybrid models (Appendix B).
This simulation-grounded approach supports direct supervision for both observable and unobservable targets, and—with a sufficient mechanistic model training set—enables out-of-the-box generalization to real-world data.
@@ 3.2 Key Capabilities
We aim for Simulation-Grounded Neural Networks (SGNNs) to address several core limitations of both mechanistic and data-driven modeling approaches. Our goal for the framework is to provide five key capabilities:
@@
  • Robustness to misspecification. By training on an ensemble of models spanning multiple structures, parameter regimes, and noise processes, SGNNs should develop flexible representations that generalize across regimes, even when real-world dynamics deviate from any one simulator.
@@
  • Mechanistic interpretability. Through back-to-simulation attribution, SGNNs will identify which simulated regimes they internally associate with real-world inputs, offering insight into the underlying dynamics the model believes are active, and enabling process-level explanations.
@@ 4.1 Overview of Evaluation Strategy
We evaluate Simulation-Grounded Neural Networks (SGNNs) across forecasting, inference, regression, and classification tasks spanning multiple scientific domains. Our experiments are designed to test the key capabilities of the SGNN framework.
4.2 SGNNs Enable Robust Forecasting Across Domains
@@ 4.2.1 SGNNs Outperform CDC COVID-19 Forecasting Models
We assessed performance using forecasting skill [28, 25, 15], defined as the percentage improvement in mean absolute error (MAE) relative to a naive baseline that predicts the most recent value:
@@
where the baseline prediction is $\hat{y}_{t+h}=y_{t}$ for all forecast horizons $h$. Normalizing by a baseline provides a dimensionless measure of skill, allowing us to compare and aggregate model performance across different locations and scales where absolute error magnitudes vary significantly. This is a stringent metric in infectious disease forecasting. The mean skill across all models in the CDC Forecast Hub was negative (–1.0) [8]. For comparison, we benchmarked the SGNN against both the median skill (13.0) and best model’s skill (34.0) of Forecast Hub models. We evaluate a full range of typical performance metrics for infectious disease performance, including weighted interval score and coverage, in appendix E.
The SGNN outperformed the baselines (Figure 2). Compared to Chronos [3], a state-of-the-art foundation model pretrained on generic time series, the SGNN achieved a roughly 12-fold improvement in skill, underscoring the importance of mechanistic grounding. The SGNN also nearly tripled the Forecast Hub’s median skill and slightly exceeded its best model performance. Finally, the SGNN outperformed a physics-informed neural network (PINN) baseline based on SEIR equations (skill = –112), which performed poorly in our tests in the presence of noise and data sparsity, consistent with prior observations [36]. This demonstrates that incorporating domain knowledge into the data rather than constraining the output of a neural network to match specific equations may sometimes be a better strategy in the presence of large uncertainty and minimal real-world data.
Figure 2: SGNNs outperform state-of-the-art baselines in real-world disease forecasting and reveal the importance of mechanistic grounding. (A) SGNNs achieve 35.3% forecasting skill on early COVID-19 mortality, almost tripling the CDC Forecast Hub median and exceeding its best model—despite using no real COVID-19 data. (B) SGNNs produce accurate forecasts with calibrated uncertainty across diverse real-world locations, including Michigan, New York, Texas, and California (four states chosen to represent US regions: West Coast, East Coast, South, and Midwest). (C) Across all U.S. states, SGNNs consistently achieve lower error (lighter) than Chronos and PINNs, showing robust generalization from synthetic pretraining. (D) On dengue, a domain with fundamentally different transmission dynamics, SGNNs outperform both PINNs and statistical models in a fully zero-shot setting—demonstrating robustness to mechanistic misspecification. (E) Replacing mechanistic simulations with neural simulators causes performance collapse, highlighting the necessity of mechanistic fidelity. (F) Restricting pretraining to simple SEIR models leads to overconfident exponential forecasts, confirming that mechanistic diversity and surveillance realism are essential for robust downstream performance.
@@ 4.2.2 SGNNs Forecast Accurately Despite Mechanistic Misspecification
Mechanistic models, whether traditional compartmental models or hybrid approaches like PINNs, rely on structural assumptions that may not hold in practice. Real-world systems often deviate from these assumptions, leading to brittle forecasts when the model form is even slightly wrong. This challenge is especially acute when key mechanisms (e.g., transmission routes or observation error) are misrepresented or omitted. To assess SGNNs under such conditions, we tested their forecasting performance with different mechanisms from those seen during pretraining.
@@
These findings underscore a strength of SGNNs: by training on a broad ensemble of mechanistic simulations spanning multiple model structures and parameter regimes, SGNNs learn flexible, scientifically grounded representations that can generalize when structural elements differ from those seen in training. This result suggests that SGNNs may not merely memorize surface patterns, but instead internalize scientific principles that extend across mechanistic boundaries. SGNNs may be learning abstract representations that capture core features of disease spread—such as delays, feedback, and depletion—that hold across different mechanistic forms.
On higher-dimensional forecasting, SGNNs maintained robust forecasting skill as the number of butterfly species increased from 4 to 16 (Figure 3A, right), outperforming task-specific neural networks (a neural network with the same architecture as the SGNN, but only trained on the historical data), which failed entirely in high dimensions. By 16 species, SGNN retained positive forecasting skill above 19.0, while baseline neural nets fell below zero. This illustrates how SGNNs can enable high-dimensional forecasting where standard approaches break down.